Sherri divided a mixed number by 3 and 3/4 . The quotient was a whole number larger than 10. What are two possibilities for the mixed number?
Answer by Alan3354(69443)

Sherri divided a mixed number by 3 and 3/4 . The quotient was a whole number larger than 10. What are two possibilities for the mixed number?
--------
Any integer multiple = 3.75*4
= n*15, n = any positive integer.